# Bellman Optimality of Average-Reward Robust Markov Decision Processes with a Constant Gain

### v1

@@ 4.1 General Criteria
Given arbitrary $\mathcal{Q}\subset\mathcal{P}(\mathcal{A})$ and $\left\{{\mathcal{P}_{s}\subset\left\{{A\rightarrow\mathcal{P}(\mathcal{S})}\right\}:s\in S}\right\}$, the following statements are equivalent:
@@
    The solutions $\left\{{v^{*}_{\gamma}:\gamma\in(0,1)}\right\}$ to the $\gamma$-discounted equation (3.3) have uniformly bounded span; i.e. $\sup_{\gamma\in(0,1)}\left|v_{\gamma}^{*}\right|_{\mathrm{span}}=\sup_{\gamma\in(0,1)}\left[\max_{s\in S}v_{\gamma}^{*}(s)-\min_{s\in S}v^{*}_{\gamma}(s)\right]<\infty$.
@@
4.2 Communicating Structures
In this section, we show that under compactness and one-sided communicating structures, the Bellman equation does have solutions.
@@ 4.2 Communicating Structures
For $p\in\mathcal{P}$ and $\Delta:S\rightarrow\mathcal{P}(\mathcal{A})$, denote $p_{\Delta}(s^{\prime}|s):=\sum_{a\in A}p(s^{\prime}|s,a)\Delta(a|s).$ Also, let $p_{\Delta}^{n}(s^{\prime}|s)$ be the $(s,s^{\prime})$ entry of the n’th power of the matrix $\left\{{p_{\Delta}(s^{\prime}|s):s,s^{\prime}\in S}\right\}$.
Definition 3.
@@
Consider arbitrary controller and adversary action sets $\mathcal{Q}\subset\mathcal{P}(\mathcal{A})$ and $\mathcal{P}=\bigtimes_{s\in S}\mathcal{P}_{s}$, with $\mathcal{P}_{s}\subset\left\{{A\rightarrow\mathcal{P}(\mathcal{S})}\right\}$.
@@
  • A stationary controller policy $\Delta:S\rightarrow\mathcal{Q}$ is said to be communicating if for any $s,s^{\prime}\in S$, there exists $p\in\mathcal{P}$ and $N\geq 1$ s.t. $p_{\Delta}^{N}(s^{\prime}|s)>0$. The controller is said to be communicating if all the stationary controller policies are communicating.
@@
  • A stationary adversary policy $p\in\mathcal{P}$ is said to be communicating if for any $s,s^{\prime}\in S$, there exists $\Delta:S\rightarrow\mathcal{Q}$ and $N\geq 1$ s.t. $p_{\Delta}^{N}(s^{\prime}|s)>0$. The adversary is said to be communicating if all the stationary adversary policies are communicating.

### v2

@@ 3.2 Bellman Optimality
We show that a solution to (3.1), if exists, will characterize the optimal robust value.
@@
In particular, the stationary policy class $\Pi_{\mathrm{S}}$ attains the same optimal value as the fully history-dependent class when playing against a history-dependent adversary whose policies belong to $\mathrm{K}_{\mathrm{H}}$; that is, $\overline{\alpha}(\mu,\Pi_{\mathrm{H}},\mathrm{K}_{\mathrm{H}})=\overline{\alpha}(\mu,\Pi_{\mathrm{S}},\mathrm{K}_{\mathrm{H}})$ and $\underline{\alpha}(\mu,\Pi_{\mathrm{H}},\mathrm{K}_{\mathrm{H}})=\underline{\alpha}(\mu,\Pi_{\mathrm{S}},\mathrm{K}_{\mathrm{H}})$. This certifies the optimality of the stationary Markov policies.
Remark 1.
Note that the $\Pi_{\mathrm{H}}$$\mathrm{K}_{\mathrm{S}}$ case is intentionally excluded from Theorem 1. This setting leads to a quite curious phenomenon, which we treat separately in Section 6. Interested readers may wish to proceed directly to that section.
Next, we show that if a solution to (3.1) also satisfies (3.2), then strong duality holds. In particular, if, in addition, the supremum and infimum are attained, the resulting policy pair constitutes a Nash equilibrium.
@@
3.3 Optimality of Stationary Policies From the Robust Bellman Equation
Similar to the classical MDP setting, stationary policy extracted from the robust Bellman equation are $\epsilon$-optimal for any $\epsilon>0$.
Theorem 3.
Let $(u^{*},\alpha^{*})$ be a solution to (3.1). If for some $\epsilon\geq 0$ and $\Delta:S\rightarrow\mathcal{Q}$,
$u^{*}(s)\leq\inf_{p_{s}\in\mathcal{P}_{s}}E_{\Delta(\cdot|s),p_{s}}[r(s,A_{0})-\alpha^{*}+u^{*}(X_{1})]+\epsilon,\quad\forall s\in S.$
Then, $\Delta$ is $\epsilon$-optimal among all stationary polices; i.e.
$\overline{\alpha}(\mu,\Pi_{\mathrm{S}},\mathrm{K})-\inf_{\kappa\in\mathrm{K}}\underline{\alpha}(\mu,\Delta,\kappa)\leq\epsilon$
where $\mathrm{K}=\mathrm{K}_{\mathrm{H}}$ or $\mathrm{K}_{\mathrm{S}}$.
Remark 2.
If we couple Theorem 3 with Theorem 1 and 2, $\Delta$ is also $\epsilon$-optimal for the robust control problem with a history-dependent controller. Moreover, note that if $\epsilon=0$, i.e. $\Delta(\cdot|s)$ achieves the $\sup_{\phi\in\mathcal{Q}}$ for all $s\in S$, then $\Delta$ is robust optimal.
@@ 4 Existence of Solution
The previous section showed that, when a solution exists, the constant-gain robust Bellman equation (3.1) characterizes the optimal value and a stationary policy. However, there is no reason to expect a constant-gain solution to exist in general, since even multichain Markov reward processes can exhibit state-dependent gain.
In this section, we develop sufficient conditions–generalizing classical MDP counterparts and motivated by operations research application–that guarantee the existence of a solution.
@@ 4.1 General Criteria
We begin with a necessary and sufficient condition that links the constant-gain average-reward equation to the discounted version. This result serves as a cornerstone for the subsequent developments.
@@
Given arbitrary $\mathcal{Q}\subseteq\mathcal{P}(\mathcal{A})$ and $\left\{{\mathcal{P}_{s}\subseteq\left\{{A\rightarrow\mathcal{P}(\mathcal{S})}\right\}:s\in S}\right\}$, the following statements are equivalent:
@@
    The solutions $\left\{{v^{*}_{\gamma}:\gamma\in(0,1)}\right\}$ to the $\gamma$-discounted equation (3.3) have uniformly bounded span; i.e.
    $\sup_{\gamma\in(0,1)}\left|v_{\gamma}^{*}\right|_{\mathrm{span}}=\sup_{\gamma\in(0,1)}\left[\max_{s\in S}v_{\gamma}^{*}(s)-\min_{s\in S}v^{*}_{\gamma}(s)\right]<\infty.$
@@
4.2 Weakly Communicating Structures
In this section, we establish that under compactness and the one-sided weakly communicating structures in Definition 3, both the robust Bellman equation (3.1) and the inf–sup equation (3.2) admit solutions. These one-sided weak-communication assumptions are well motivated: they are extensively used in the classical MDP literature to guarantee a constant optimal gain [18]. In particular, as reviewed earlier, one-sided communication ensures the existence of solutions in SG settings where both players may employ fully randomized strategies [27], which in turn induces convex controller policy and adversary ambiguity sets. However, to the best of our knowledge, extensions to the weak-communication setting have not been established in the literature.
Moreover, the robust MDP landscape is more intricate, as applications sometimes necessitate non-convex decision sets for one or both players. We show that, under one-sided weak communication and compactness—without requiring convexity—the constant-gain robust Bellman equation (3.1) or its inf–sup counterpart (3.2) admits a solution, with the relevant equation determined by which side satisfies the weak-communication assumption.
We proceed by introducing the following notation. For $p\in\mathcal{P}$ and $\Delta:S\rightarrow\mathcal{P}(\mathcal{A})$, denote
$p_{\Delta}(s^{\prime}|s):=\sum_{a\in A}p(s^{\prime}|s,a)\Delta(a|s).$
@@ 4.2 Weakly Communicating Structures
Also, let $p_{\Delta}^{n}(s^{\prime}|s)$ be the $(s,s^{\prime})$ entry of the n’th power of the matrix $\left\{{p_{\Delta}(s^{\prime}|s):s,s^{\prime}\in S}\right\}$. Moreover, for $C\subseteq S$ we denote the complement of $C$ is $S$ by $C^{c}:=S\setminus C$.
Definition 3 (Weak Communication).
@@
Consider arbitrary controller and adversary action sets $\mathcal{Q}\subseteq\mathcal{P}(\mathcal{A})$ and $\mathcal{P}=\bigtimes_{s\in S}\mathcal{P}_{s}$, with $\mathcal{P}_{s}\subseteq\left\{{A\rightarrow\mathcal{P}(\mathcal{S})}\right\}$.
@@
  • A stationary controller policy $\Delta:S\rightarrow\mathcal{Q}$ is said to be weakly communicating if there is a communicating class $C_{\Delta}\subseteq S$ s.t. for any $s,s^{\prime}\in C_{\Delta}$, there exists $p\in\mathcal{P}$ and $N\geq 1$ s.t. $p_{\Delta}^{N}(s^{\prime}|s)>0$. Moreover, for all $\in C_{\Delta}^{c}, ss$ is transient under any stationary adversarial policy.
@@
    The controller is weakly communicating if every stationary policy $\Delta:S\rightarrow\mathcal{Q}$ is weakly communicating.
  • A stationary adversary policy $p\in\mathcal{P}$ is said to be weakly communicating if there is a communicating class $C_{p}\subseteq S$ s.t. for any $s,s^{\prime}\in C_{p}$, there exists $\Delta:S\rightarrow\mathcal{Q}$ and $N\geq 1$ s.t. $p_{\Delta}^{N}(s^{\prime}|s)>0$. Moreover, for all $\in C_{p}^{c}, ss$ is transient under any stationary controller policy.
    The adversary is weakly communicating if every stationary policy $p\in\mathcal{P}$ is weakly communicating.
Communicating controller and adversary are defined analogously as follows.
Definition 4 (Communication).
Consider arbitrary controller and adversary action sets $\mathcal{Q}\subseteq\mathcal{P}(\mathcal{A})$ and $\mathcal{P}=\bigtimes_{s\in S}\mathcal{P}_{s}$, with $\mathcal{P}_{s}\subseteq\left\{{A\rightarrow\mathcal{P}(\mathcal{S})}\right\}$.
  • A stationary controller policy $\Delta:S\rightarrow\mathcal{Q}$ is communicating if it is weakly communicating with $C_{\Delta}=S$.
    The controller is communicating if every stationary policy $\Delta:S\rightarrow\mathcal{Q}$ is communicating.
  • A stationary adversary policy $p\in\mathcal{P}$ is communicating if it is weakly communicating with $C_{p}=S$.
    The adversary is communicating if every stationary policy $p\in\mathcal{P}$ is communicating.
Remark.
Note that our (weak) communication definitions parallel their classical MDP counterpart. Moreover, a controller/adversary may be (weakly) communicating even when the communicating sets $C_{\Delta}$$C_{p}$ depend on the particular stationary policy; they need not coincide across stationary policies.
With these definitions, we are ready to state the main results of this section.
Theorem 5 (Controller-Side Structures).
Assume either of the two assumptions holds:
  1. (1)
    The controller is weakly communicating and $\mathcal{Q}$ and $\mathcal{P}_{s}$ are compact for all $s\in S$.
  2. (2)
    The controller is communicating and $\mathcal{Q}$ is compact.
Then the constant gain average reward robust Bellman equation (3.1) has a solution.
The proof of Theorem 5 is provided within the main body of the paper in Section 5.
Symmetric to the controller-side results, we show that the adversary-side weak communication structures will imply the existence of solutions to (3.2). The proof for Theorem 6 is deferred to Appendix D.
Theorem 6 (Adversary-Side Structures).
Assume either of the two assumptions holds:
  1. (1)
    The adversary is weakly communicating and $\mathcal{Q}$ and $\mathcal{P}_{s}$ are compact for all $s\in S$.
  2. (2)
    The adversary is communicating and $\mathcal{P}_{s}$ is compact for all $s\in S$.
Then the constant gain average reward inf-sup equation (3.2) has a solution.
Note that Theorem 6 does not, by itself, guarantee a solution to (3.1). Nevertheless, under suitable sufficient conditions, the solution $(u^{\prime},\alpha^{\prime})$ of (3.2) also satisfies (3.1). Let $\delta_{a}\in\mathcal{P}(\mathcal{A})$ denote the Dirac measure at $a\in A$.
@@
Either of the following conditions is sufficient for a solution $(u^{\prime},\alpha^{\prime})$ of (3.2) to also solve (3.1):
  1. (1)
    $\mathcal{Q}$ and each $\mathcal{P}_{s}$ are convex for all $s\in S$, and either $\mathcal{Q}$ is compact or all $\mathcal{P}_{s},s\in S$ are compact.
  2. (2)
    $\left\{{\delta_{a}:a\in A}\right\}\subseteq\mathcal{Q}$, i.e., the controller can take deterministic actions, and for every $s\in S$, the ambiguity set factorizes as $\mathcal{P}_{s}=\bigtimes_{a\in A}\mathcal{P}_{s,a}$ for some $\mathcal{P}_{s,a}\subseteq\mathcal{P}(S)$. In this case, we refer to $\mathcal{P}$ as an SA-rectangular adversarial ambiguity set.
4.3 Useful Corollaries
Although the conditions of Theorems 5 and 6 are self-explanatory, and their verification mirrors that of their non-robust counterparts, in this section we highlight several scenarios, motivated by operations applications, in which Theorems 5 and 6 apply.
Stability structures are ubiquitous in operations research applications: reasonable policies typically lead to systems that are stable and insensitive to their initial states. We show that, under stability assumptions such as (weak) irreducibility or (strong) unichain, Theorems 5 and 6 are verified.
Corollary 7.1 (Irreducible).